\begin{thebibliography}{10}

\bibitem{Chaundy}Burchnall J.-L., Chaundy T.W. Commutative ordinary differential operators.
Proc. London Math. Soc. 21 (1923), 420-440; Proc. Royal Soc. London (A)
118 (1928), 557-583.
\bibitem{theta}I. M. Krichever, Integration of nonlinear equations by the methods of algebraic geometry, Func-
tional Anal. Appl., 11: 1 (1977), 12–26.
\bibitem{ringkrichever}I.M. Krichever. Commutative rings of ordinary linear differential operators.Funktsional. Anal. i Prilozhen.,12:3 (1978), 20–31 (in Russian);English translation inFunctional Anal. Appl.,12:3 (1978), 175–185.
\bibitem{novkrich}I. M. Krichever, S.P. Novikov. Holomorphic bundles and nonlinear equations. Finite-zone solutions of rank 2.Dokl. Akad. Nauk SSSR,247:1 (1979), 33–36(in Russian);English translation in Soviet Math. Dokl..
\bibitem{Mokhov1}O. I. Mokhov, Commuting ordinary differential operators of rank 3 corresponding to an elliptic curve. Uspekhi Matem. Nauk,37:4 (1982), 169–170 (in Russian); English translation in Russian Math. Surveys,37:4 (1982), 129–130.
\bibitem{Mokhov2}O. I. Mokhov, Commuting differential operators of rank 3, and nonlinear differential equations.Izvestiya AN SSSR, ser. matem.,53:6 (1989), 1291–1315 (in Russian); English translation in Math. USSR, Izvestiya,35:3 (1990), 629–655.
\bibitem{Mironov} A. E. Mironov, Self-adjoint commuting differential operators and commutative subalgebras of the Weyl algebra. Invent. math. (2014) 197:417-431
\bibitem{Dixmier} J. Dixmier, Bull. Soc. Math. France,96(1968), 209–24
\bibitem{Mironov2} A. E. Mironov, A.E. Mironov.  Periodic and rapid decay rank two self-adjoint commuting differential operators, Amer. Math. Soc. Transl. Ser. 2, V. 234, 2014, P. 309--322.
\bibitem{Mokhov3} O. I. Mokhov. On Commutative Subalgebras of the Weyl Algebra Related to Commuting Operators of Arbitrary Rank and Genus”, Mat. Zametki, 94:2 (2013), 314–316 (in Russian); English translation in Mathematical Notes, vol. 94:2 (2013), 298-300
\bibitem{Mokhov4}  O. I. Mokhov, Commuting ordinary differential operators of arbitrary genus and arbitrary rank with polynomial coefficients. American Mathematical Society Translations, Volume 234 (2014),  323-336.
\bibitem{Smirnov} A. O. Smirnov, Finite-gap elliptic solutions of the KdV equation. Acta Appl. Math., 36 (1994), 125-166.
\bibitem{Smirnov4} A. O. Smirnov, Real finite-gap regular solutions of the Kaup-Boussinesq equation,
Theoret. and Math. Phys., 66 (1986), 19-31.
\bibitem{Vartan}  V. Oganesyan, Commuting differential operators of rank 2 with polynomial coefficients, Functional Analysis and Its Applications , Vol . 50, No . 1,  arXiv:1409.4058.
\bibitem{Ince}  E. L. Ince, Ordinary differential equations, Dover, New York, 1956.
\bibitem{Wal}  G. Wallenberg, Uber die Vertauschbarkeit homogener linearer Differentialausdrucke. Arch.
Math. Phys. 4 (1903), 252–268.
\bibitem{Schur} J. Schur, Uber vertauschbare lineare Differentialausdrucke. Sitzungsber. der Berliner Math.
Gesell. 4 (1905), 2–8.
\bibitem{Vartan2} V. Oganesyan,  Explicit characterization of some commuting differential operators of rank 2, arXiv:1502.07491.
\bibitem{Vartan3} V. Oganesyan,  "Commuting differential operators of rank 2 and arbitrary genus g with polynomial coefficients", Russian Math. Surveys, 2015, 70 (1), 165--167.
\bibitem{Davl} V. N. Davletshina, Siberian Mathematical Journal May 2015, Volume 56, Issue 3, pp 405-410.
\bibitem{Zegl}  Andrey E. Mironov, Alexander B. Zheglov "Commuting Ordinary Differential Operators with Polynomial Coefficients and Automorphisms of the First Weyl Algebra", Int Math Res Notices 2015 : rnv218v1-20.
\bibitem{Grin} P. G. Grinevich, "Vector rank of commuting matrix differential operators. Proof of S. P. Novikov's criterion", Math USSR IZV, 1987, 28 (3), 445–465.
\bibitem{weikard1} F. Gesztesy , K. Unterkofler , R. Weikard, An explicit characterization of Calogero–Moser systems, Trans. Amer. Math. Soc.
\bibitem{weikard2} R. Weikard, On commuting matrix differential operators, New York Journal of Mathematics.
\bibitem{Dub} B.~A.~Dubrovin, V.~B.~Matveev, S.~P.~Novikov, Non-linear equations of Korteweg--de Vries type, finite-zone linear
operators, and Abelian varieties,  Uspekhi Mat. Nauk, 1979, 31, 59--146.
\bibitem{weikard3} F. Gesztesy, R. Weikard,  A characterization of all elliptic algebro-geometric solutions of the AKNS hierarchy, Acta Mathematica, September 1998, Volume 181, Issue 1, pp 63-108.
\bibitem{Dubrovin1} B. A. Dubrovin, “Completely integrable Hamiltonian systems associated with matrix operators and Abelian varieties”, Funktsional. Anal. i Prilozhen., 11:4 (1977), 28–41.
\bibitem{Mumford} D. Mumford,  An algebro-geometric construction of commuting operators and of solutions to the Toda lattice equation, Korteweg deVries equation and related nonlinear equation. Proceedings of the International Symposium on Algebraic Geometry (Kyoto Univ., Kyoto, 1977), pp. 115–153, Kinokuniya Book Store, Tokyo, 1978.
\bibitem{Dickey} L. A. Dickey, On the $\tau$-function of matrix hierarchies of integrable equations. J. Math. Phys. 32 (1991), no. 11, 2996–3002.
\bibitem{Levinson} E. A. Coddington, N. Levinson, Theory of Ordinary Differential Equations. Krieger, Malabar, 1984.
\end{thebibliography}

\documentclass[13pt,a4paper]{article}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{euscript}
\usepackage[cp1251]{inputenc}
\usepackage[english] {babel}
\usepackage[pdftex]{graphicx}
\usepackage[12pt]{extsizes}

\begin{document}
\let\thefootnote\relax\footnote{This work was supported in part by the Simons Foundation.}
\begin{center}
  \textbf{AKNS hierarchy and Schrodinger finite-gap  potentials}
\end{center}
\begin{center}
  \textbf{V. Oganesyan}
\end{center}
\begin{center}
   \textbf{1. Introduction}
\end{center}

The commutativity condition $L_nL_m = L_mL_n$ of two differential operators
\begin{equation*}
L_n= \sum\limits^{n}_{i=0} u_i(x)\partial_x^i,  \quad  L_m= \sum\limits^{m}_{i=0} v_i(x)\partial_x^i,
\end{equation*}
where coefficients $u_i(x)$ and $v_i(x)$ are scalar or matrix valued functions, is equivalent to a very complicated system of nonlinear differential equations. The theory of commuting ordinary differential operators was first developed in the beginning of the XX century in the works of Wallenberg ~\cite{Wal}, Schur ~\cite{Schur}, and Burchnall, Chaundy ~\cite{Chaundy}. \\
If two differential operators with scalar or matrix valued coefficients commute, then there exists a nonzero polynomial $R(z,w)$ such that  $R(L_n,L_m)=0$ (see ~\cite{Chaundy}, ~\cite{Grin}). The curve $\Gamma$ defined by $R(z,w)=0$ is called the \emph{spectral curve} of this pair of operators.\\
If coefficients are scalar functions and
\begin{equation*}
L_n \psi=z\psi, \quad  L_m \psi=w\psi,
\end{equation*}
then $(z,w) \in \Gamma$. For almost all $(z,w) \in \Gamma$, the dimension of the space of common eigenfunctions $\psi$ is the same. The dimension of the space of common eigenfunctions of two commuting scalar differential operators is called the \emph{rank} of this pair. The rank is a common divisor of m and n. The genus of the spectral curve of a pair of commuting operators is sometimes called the genus
of this pair.\\
When coefficients are matrix of size $s\times s$ we have the \emph{vector rank} $(l_1,...,l_k)$, where $k\leqslant s$ (see ~\cite{Grin}). Numbers $l_i$ are common divisors of $m$ and $n$. \\
The first examples of pairs of commuting  scalar differential operators of the nontrivial ranks 2 and 3 and the nontrivial genus g=1 were constructed by Dixmier ~\cite{Dixmier} for the nonsingular elliptic spectral curve $w^2=z^3-\alpha$, where $\alpha$ is and arbitrary nonzero constant.\\
A general classification of commuting scalar  differential operators was obtained by Krichever ~\cite{ringkrichever}. Unfortunately this classification helps to find coefficients of commuting operators explicitly only when rank equals $1$ and in some other cases. The general form of commuting operators of rank 2 for an arbitrary elliptic spectral curve was found by Krichever and Novikov ~\cite{novkrich}. The general form of operators of rank 3 with arbitrary elliptic spectral curve was found by Mokhov ~\cite{Mokhov1},~\cite{Mokhov2}. Mironov in ~\cite{Mironov} found new methods of constructing new scalar commuting operators of rank 2. Using Mironov's method many examples of scalar commuting operators of rank 2 were found (see ~\cite{Mokhov4}, ~\cite{Vartan3} ~\cite{Zegl}, ~\cite{Vartan}, ~\cite{Vartan2}, ~\cite{Davl}).\\
A general classification of commuting matrix   differential operators  was obtained by Grinevich ~\cite{Grin}.\\
Many nonlinear equations are equivalent to commutativity condition of scalar or matrix differential operators. Using theory of commuting operators many exact solutions of nonlinear differential equations were found (see ~\cite{Smirnov},  ~\cite{Smirnov4}, ~\cite{Dubrovin1}, ~\cite{Mumford}, ~\cite{Dickey}).
\\
\\
The author is grateful to Professor O. I. Mokhov and Professor A. O. Smirnov for valuable discussions.
\begin{center}
\textbf{2. AKNS hierarchy and finite-gap Schrodinger potentials}
\end{center}

Let us consider
\begin{equation*}
M= \begin{pmatrix}
\imath & 0  \\
0 & -\imath  \\
\end{pmatrix}
\partial_x +
 \begin{pmatrix}
0 & -\imath q(x)  \\
\imath p(x) & 0  \\
\end{pmatrix}.
\end{equation*}
Denote by $L_n$ any differential operator of order $n$  with matrix coefficients of size $2\times 2$.\\
Define functions $f_k$, $g_k$ by the recurrence relations
\begin{equation}
 \begin{cases}
   f_{k+1}=\frac{i}{2}f'_{k}(x) - iq(x)(\int(p(x)f_{k}(x) + q(x)g_{k}(x))dx + C_k)\\
   g_{k+1}=-\frac{i}{2}g'_{k}(x) + ip(x)(\int(p(x)f_k(x) + q(x)g_k(x))dx + C_k)\\
 \end{cases}
\end{equation}
where $C_k$ are arbitrary constants and
\begin{equation*}
 \begin{cases}
   f_{1}=-iq(x)\\
   g_{1}=ip(x)\\
 \end{cases}
\end{equation*}
For example
\begin{equation*}
 \begin{cases}
   f_{3}= \frac{i}{4}q_{xx} - \frac{i}{2}pq^2 + \frac{C_1}{2}q_x - iC_2q\\
   g_{3}= - \frac{i}{4}p_{xx} + \frac{i}{2}p^2q + \frac{C_1}{2}p_x + iC_2p.\\
 \end{cases}
\end{equation*}
Easy to check that $f_k$ and $g_k$ are polynomials in $p,q,p',q',...p^{(k-1)},q^{(k-1)}$ (see \cite{weikard3}).  Functions $p(x), q(x)$ are solutions of
\begin{equation}
 \begin{cases}
   f_{m+1}=0\\
   g_{m+1}=0\\
 \end{cases}
\end{equation}
if and only if there exists an operator $L_{m+1}$ such that $ML_{m+1}=L_{m+1}M$ (see ~\cite{weikard3}) and
\begin{equation*}
L_{m+1}^2 = M^{2m+2} + b_{2m}M^{2m} + ... + b_1M + b_{0},
\end{equation*}
where $b_i$ are certain constants.
Easy to see from (1) that if condition (2) is satisfied, then
\begin{equation}
 \begin{cases}
   f_{k}=0\\
   g_{k}=0\\
 \end{cases}
\end{equation}
for all $k \geqslant m+1$.  \\
Assume that $p(x)$ and $q(x)$ have isolated poles of order $1$ at points $a_i$. And in a neighborhood of $a_i$ the Laurent series have the form
\begin{equation*}
p(x)=\frac{\varphi_{i,-1}}{x-a_i} + \varphi_{i,0} + \varphi_{i,1}(x-a_i)+...
\end{equation*}
\begin{equation*}
q(x)=\frac{\psi_{i,-1}}{x-a_i} + \psi_{i,0} + \psi_{i,1}(x-a_i)+...
\end{equation*}
The main results of this paper are the following.\\
\textbf{Theorem 1.}\\
\emph{If functions $p(x), q(x)$ are solutions of equations $f_{2m+2}=0$ and $g_{2m+2}=0$ or  $f_{2m+3}=0$ and $g_{2m+3}=0$, then}
\begin{equation*}
\begin{gathered}
\varphi_{i,-1} = \frac{n^2_i}{\psi_{i,-1}}, \varphi_{i,0} = -\frac{\psi_{i,0}\varphi_{i,-1}^2}{n^2_i},  \varphi_{i,1} = \frac{\psi_{i,1}\varphi_{i,-1}^2}{n^2_i}, \varphi_{i,2} = -\frac{\psi_{i,2}\varphi_{i,-1}^2}{n^2_i},\\ \varphi_{i,3} = \frac{\psi_{i,3}\varphi_{i,-1}^2}{n^2_i},..., \varphi_{i,2n_i-1} = \frac{\psi_{i,2n_i-1}\varphi_{i,-1}^2}{n^2_i},\\
\end{gathered}
\end{equation*}
\emph{where $n_i \in \mathbb{N}$ and $n_i \leqslant m$.}\\
\textbf{Theorem 2.}\\
\emph{If}
\begin{equation*}
\begin{gathered}
\varphi_{i,-1} = \frac{n^2_i}{\psi_{i,-1}}, \varphi_{i,0} = -\frac{\psi_{i,0}\varphi_{i,-1}^2}{n^2_i},  \varphi_{i,1} = \frac{\psi_{i,1}\varphi_{i,-1}^2}{n^2_i}, \varphi_{i,2} = -\frac{\psi_{i,2}\varphi_{i,-1}^2}{n^2_i},\\ \varphi_{i,3} = \frac{\psi_{i,3}\varphi_{i,-1}^2}{n^2_i},..., \varphi_{i,2n_i-1} = \frac{\psi_{i,2n_i-1}\varphi_{i,-1}^2}{n^2_i},\\
\end{gathered}
\end{equation*}
for some $n_i$, then solutions of
\begin{equation*}
\begin{pmatrix}
\imath & 0  \\
0 & -\imath  \\
\end{pmatrix}
\partial_xy +
 \begin{pmatrix}
0 & -\imath q(x)  \\
\imath p(x) & 0  \\
\end{pmatrix}y = zy
\end{equation*}
\emph{are meromorphic in a neighborhood of a pole $a_i$ for all  $z \in \mathbb{C}$, where $y=(y_1(x),y_2(x))^{t}$.}\\
\\
We know the following theorem (see ~\cite{weikard2})\\
\\
\textbf{Theorem.}\\
\emph{Assume that $p(x)$, $q(x)$ are rational functions bounded at infinity or else that $p(x)$, $q(x)$ are  meromorphic $\omega$-periodic functions with finitely many poles in the period strip. Then functions $p(x)$, $q(x)$ are solutions of some equations of AKNS hierarchy  if and only if all solutions of
\begin{equation*}
\begin{pmatrix}
\imath & 0  \\
0 & -\imath  \\
\end{pmatrix}
\partial_xy +
 \begin{pmatrix}
0 & -\imath q(x)  \\
\imath p(x) & 0  \\
\end{pmatrix}y = zy
\end{equation*}
are meromorphic for any  $z \in \mathbb{C}$, where $y=(y_1(x),y_2(x))^{t}$.}\\
\\
So, we obtain the following Theorem\\
\\
\textbf{Theorem 3.}\\
\emph{Assume that $p(x)$, $q(x)$ are rational functions bounded at infinity or else that $p(x)$, $q(x)$ are  meromorphic $\omega$-periodic functions with finitely many poles in the period strip. Assume that all poles of $p(x)$ and $q(x)$ are poles of order $1$. Then functions $p(x)$ and $q(x)$ are solutions of $f_{m}=0$ and $g_{m}=0$ if and only if}
\begin{equation*}
\begin{gathered}
\varphi_{i,-1} = \frac{n^2_i}{\psi_{i,-1}}, \varphi_{i,0} = -\frac{\psi_{i,0}\varphi_{i,-1}^2}{n^2_i},  \varphi_{i,1} = \frac{\psi_{i,1}\varphi_{i,-1}^2}{n^2_i}, \varphi_{i,2} = -\frac{\psi_{i,2}\varphi_{i,-1}^2}{n^2_i},\\ \varphi_{i,3} = \frac{\psi_{i,3}\varphi_{i,-1}^2}{n^2_i},..., \varphi_{i,2n_i-1} = \frac{\psi_{i,2n_i-1}\varphi_{i,-1}^2}{n^2_i}.\\
\end{gathered}
\end{equation*}
for some $n_i \in \mathbb{N}$.
\\
\\
Assume that $p(x)$, $q(x)$ satisfy the condition of Theorem 1. Let us consider the function $pq$. The Laurent series of function $pq$ in a neighborhood of $a_i$ has the form
\begin{equation*}
pq= \frac{h_{i,-2}}{(x-a_i)^2} + \frac{h_{i,-1}}{(x-a_i)} + h_{i,0} + h_{i,1}(x-a_i) + O((x-a_i)^2),
\end{equation*}
where
\begin{equation*}
\begin{gathered}
h_{i,-2} =n^2_i,\\
h_{i,2k-1} = (\varphi_{i,-1}\psi_{i,2k} + \psi_{i,-1}\varphi_{i, 2k}) + (\varphi_{i,0}\psi_{i,2k-1} + \psi_{i,0}\varphi_{i, 2k-1}) +...+\\
+(\varphi_{i,l}\psi_{i,2k-l-1} + \psi_{i,l}\varphi_{i, 2k-l-1}) + ...+(\varphi_{i,2k}\psi_{i,-1} + \psi_{i,2k}\varphi_{i, -1}).
\end{gathered}
\end{equation*}
But we know that $p$ and $q$ satisfy the condition of Theorem 1 and
\begin{equation*}
\varphi_{i,l}\psi_{i,2k-l-1} + \psi_{i,l}\varphi_{i, 2k-l-1} = \varphi_{i,-1}\psi_{i,-1}(\frac{\varphi_{i,l}\psi_{i,2k-l-1}}{\varphi_{i,-1}\psi_{i,-1}} + \frac{\psi_{i,l}\varphi_{i,2k-l-1}}{\varphi_{i,-1}\psi_{i,-1}} )=
\end{equation*}
\begin{equation*}
=\psi_{i,-1}(\frac{\psi_{i,2k-l-1}}{\psi_{i,-1}} + (-1)^{l+1}\frac{\varphi_{i,2k-l-1}}{\varphi_{i,-1}} ) = (\frac{\psi_{i,2k-l-1}}{\psi_{i,-1}} + (-1)^{2k-l-1}\frac{\varphi_{i,2k-l-1}}{\varphi_{i,-1}} ) = 0.
\end{equation*}
So, we see that coefficients $h_{i,-1} = h_{i,1}=...=h_{i,2n_i-1} =0$. We obtain that in a neighborhood of $a_i$ the Laurent series of function $pq$ have the form
\begin{equation*}
pq = \frac{n^2_i}{(x-a_i)^2} + \sum^{t=n_i}_{t=0}h_{i,2t}(x-a_i)^t +  h_{i, 2n_i+1}(x-a_i)^{2n_i+1} + O((x-a_i)^{2n_i+2})
\end{equation*}
\textbf{Theorem 4.} \emph{Let us suppose that $p(x)$ and $q(x)$ are rational or simply periodic functions with a common period. Let us assume that $p(x)$ and $q(x)$ haven't isolated singularity at infinity. Suppose that functions $p(x)$, $q(x)$ have finitely many poles in the period strip. Also suppose that all poles of $p,q$ are poles of order $1$ and $\varphi_{i,-1}\psi_{i,-1}=n^2$ for any $i$. If $p(x)$, $q(x)$ are solutions of $f_{m}=0$ and $g_{m}=0$ for any $m>2$, then function $u(x)=\frac{n+1}{n}pq$ is finite-gap Schrodinger  potential. }\\
\\
Let us consider some examples (see \cite{weikard3}, page 101)\\
\textbf{Example 1.}\\
Functions $p=\frac{\alpha}{\sin(x)}$ and $q=\frac{\beta}{\sin(x)}$, where $\alpha\beta=n^2$ and $n \in \mathbb{N}$, are solutions of $f_{2n+2}=0$ and $g_{2n+2}=0$. Product $pq=\frac{n^2}{\sin^2(x)}$  and  $\frac{n+1}{n}pq=\frac{n(n+1)}{\sin^2(x)}$ is Schrodinger finite-gap potential.\\
\textbf{Example 2.}\\
 Let us consider $p=\alpha(\zeta(x) - \zeta(x-\omega_2) - \zeta(\omega_2))$ and $q=\beta(\zeta(x) - \zeta(x-\omega_2) - \zeta(\omega_2))$, where $\zeta(x; \omega_1,\omega_2)$ is  Weierstrass function, $\alpha\beta = n^2$, $n \in \mathbb{N}$. Functions $p, q$ are solutions of  $f_{m+1}=0$ and $g_{m+1}=0$ for some $m$. Function $\frac{n+1}{n}pq=n(n+1)\wp(x) + n(n+1)\wp(x-\omega_2)$ is Schrodinger finite-gap potential.
\begin{center}
\textbf{3. Proof of Theorem 1}
\end{center}

We know that in a neighborhood of $a_i$
\begin{equation*}
\begin{gathered}
f_1=-\frac{i\psi_{i,-1}}{(x-a_i)} + O(1) = \frac{A^1_{i,-1}}{(x-a_i)} + O(1)\\
g_1=\frac{i\varphi_{i,-1}}{(x-a_i)} + O(1) = \frac{B^1_{i,-1}}{(x-a_i)} + O(1).
\end{gathered}
\end{equation*}
By $A^k_{i,m}$ denote a coefficient in the term $(x-a_i)^m$ in Laurent series  of $f_k$ at point $a_i$. Similarly by $B^k_{i,m}$ denote a coefficient in the term $(x-a_i)^m$ in Laurent series  of $g_k$ at point $a_i$.\\
\textbf{Lemma 1.} There exists $n_i\in \mathbb{N}$ such that $\varphi_{i,-1}=\frac{n^2_i}{\psi_{i,-1}}$ and $B^{k}_{i,-k} = (-1)^k\frac{\varphi_{i,-1}A^{k}_{i,-k}}{\psi_{i,-1}}$.\\
\textbf{Proof.}\\
We obviously have $B^1_{i,-1} = -\frac{\varphi_{i,-1}A^1_{i,-1}}{\psi_{i,-1}}$.\\
Let us calculate $f_2$ and $g_2$
\begin{equation*}
\begin{gathered}
f_2=-\frac{\psi_{i,-1}}{2(x-a_i)^2} + O(\frac{1}{(x-a_i)}) = \frac{A^2_{i,-2}}{(x-a_i)^2} + O(\frac{1}{(x-a_i)})\\
g_2=-\frac{\varphi_{i,-1}}{2(x-a_i)^2} + O(\frac{1}{(x-a_i)}) = \frac{B^2_{i,-2}}{(x-a_i)^2} + O(\frac{1}{(x-a_i)})
\end{gathered}
\end{equation*}
We get $B^2_{i,-2} = \frac{\varphi_{i,-1}A^2_{i,-2}}{\psi_{i,-1}}$. Consider $f_3$ and $g_3$
\begin{equation*}
\begin{gathered}
f_3=-\frac{i\psi_{i,-1}(\varphi_{i,-1}\psi_{i,-1}- 1)}{2(x-a_i)^3} + O(\frac{1}{(x-a_i)^2}) = \frac{A^3_{i,-3}}{(x-a_i)^3} + O(\frac{1}{(x-a_i)^2})\\
g_3=\frac{i\varphi_{i,-1}(\varphi_{i,-1}\psi_{i,-1} - 1)}{2(x-a_i)^3} + O(\frac{1}{(x-a_i)^2}) = \frac{B^3_{i,-3}}{(x-a_i)^3} + O(\frac{1}{(x-a_i)^2})
\end{gathered}
\end{equation*}
We get $B^3_{i,-3} = -\frac{\varphi_{i,-1}A^3_{i,-3}}{\psi_{i,-1}}$. Let us prove that
\begin{equation}
B^{2k+1}_{i,-2k-1} = -\frac{\varphi_{i,-1}A^{2k+1}_{i,-2k-1}}{\psi_{i,-1}}, \quad B^{2k+2}_{i,-2k-2} = \frac{\varphi_{i,-1}A^{2k+2}_{i,-2k-2}}{\psi_{i,-1}}.
\end{equation}
The proof is by induction on $k$. We checked this for $k=0$ and $k=1$. By the induction hypothesis,
\begin{equation*}
B^{2k+1}_{i,-2k-1} = -\frac{\varphi_{i,-1}A^{2k+1}_{i,-2k-1}}{\psi_{i,-1}}, \quad B^{2k}_{i,-2k} = \frac{\varphi_{i,-1}A^{2k}_{i,-2k}}{\psi_{i,-1}}.
\end{equation*}
We obtain
\begin{equation}
\begin{gathered}
f_{2k+2}=\frac{2i\psi_{i,-1}(A^{2k+1}_{i,-2k-1}\varphi_{i,-1} + B^{2k+1}_{i,-2k-1}\psi_{i,-1}) - iA^{2k+1}_{i,-2k-1}(2k+1)^2}{2(2k+1)(x-a_i)^{2k+2}} + O(\frac{1}{(x-a_i)^{2k+1}}) =\\
=-\frac{iA^{2k+1}_{i,-2k-1}(2k+1)}{2(x-a_i)^{2k+2}} + O(\frac{1}{(x-a_i)^{2k+1}}),
\end{gathered}
\end{equation}
\begin{equation}
\begin{gathered}
g_{2k+2}=\frac{iB^{2k+1}_{i,-2k-1}(2k+1)^2 - 2i\varphi_{i,-1}(A^{2k+1}_{i,-2k-1}\varphi_{i,-1} + B^{2k+1}_{i,-2k-1}\psi_{i,-1} ) }{2(2k+1)(x-a_i)^{2k+2}} + O(\frac{1}{(x-a_i)^{2k+1}}) =\\
=-\frac{iA^{2k+1}_{i,-2k-1}\varphi_{i,-1}(2k+1)}{2\psi_{i,-1}(x-a_i)^{2k+2}} + O(\frac{1}{(x-a_i)^{2k+1}})
\end{gathered}
\end{equation}
In this case  we see that $B^{2k+2}_{i,-2k-2} = \frac{\varphi_{i,-1}A^{2k+2}_{i,-2k-2}}{\psi_{i,-1}}$. Then
\begin{equation}
\begin{gathered}
f_{2k+3}=-i\frac{2(k+1)^2A^{2k+2}_{i,-2k-2} - \psi_{i,-1}(B^{2k+2}_{i,-2k-2}\psi_{i,-1} + A^{2k+2}_{i,-2k-2}\varphi_{i,-1})}{2(k+1)(x-a_i)^{2k+3}} + O(\frac{1}{(x-a_i)^{2k+2}}) =\\
=\frac{iA^{2k+2}_{i,-2k-2}(\varphi_{i,-1}\psi_{i,-1} - (k+1)^2)}{(k+1)(x-a_i)^{2k+3}} + O(\frac{1}{(x-a_i)^{2k+2}})
\end{gathered}
\end{equation}
\begin{equation}
\begin{gathered}
g_{2k+3}=i\frac{2(k+1)^2B^{2k+2}_{i,-2k-2} - \varphi_{i,-1}(B^{2k+2}_{i,-2k-2}\psi_{i,-1} + A^{2k+2}_{i,-2k-2}\varphi_{i,-1}) }{2(k+1)(x-a_i)^{2k+3}} + O(\frac{1}{(x-a_i)^{2k+2}}) =\\
=-\frac{iA^{2k+2}_{i,-2k-2}\varphi_{i,-1}(\varphi_{i,-1}\psi_{i,-1} - (k+1)^2)}{(k+1)\psi_{i,-1}(x-a_i)^{2k+3}} + O(\frac{1}{(x-a_i)^{2k+2}}).
\end{gathered}
\end{equation}
So (4) is proved.\\
We know that there exists $l$ such that $f_{l}=0$ and $g_{l}=0$ hence $A^l_{i,-l}=B^l_{i,-l}=0$ for some $l$. But we see from (5), (6) and (7), (8) that  $A^k_{i,-k}=B^k_{i,-k}=0$  if and only if $\varphi_{i,-1}\psi_{i,-1}=n^2_i$.\\
\textbf{The Lemma is proved.}\\
\\
\textbf{Lemma 2.} $\varphi_{i,0}=-\frac{\psi_{i,0}\varphi_{i,-1}^2}{n^2_i}$.\\
\textbf{Proof.}\\
As in the previous Lemma let us consider
\begin{equation*}
\begin{gathered}
f_1=-\frac{i\psi_{i,-1}}{(x-a_i)} -i\psi_{i,0} + O((x-a_i)) ,\\
g_1=\frac{i\varphi_{i,-1}}{(x-a_i)} + i\varphi_{i,0} + O((x-a_i)).
\end{gathered}
\end{equation*}
We see that $\frac{A^1_{i,0}}{\psi_{i,-1}} - \frac{B^1_{i,0}}{\varphi_{i,-1}} = -\frac{i}{n_i^2}(\psi_{i,-1}\varphi_{i,0} + \varphi_{i,-1}\psi_{i,0})$. Let us prove that
\begin{equation}
\frac{A^p_{i,-p+1}}{\psi_{i,-1}} + (-1)^p \frac{B^p_{i,-p+1}}{\varphi_{i,-1}} = \alpha_p(\psi_{i,-1}\varphi_{i,0} + \varphi_{i,-1}\psi_{i,0}),
\end{equation}
where $\alpha_p \neq 0$ and doesn't depend on $C_i$, $p\in \mathbb{N}$ is an arbitrary number. The proof is by induction on $p$. Suppose that the assumption is true for $A^p_{i,-p+1}$ and $B^p_{i,-p+1}$. Easy to see that
\begin{equation}
\begin{gathered}
A^{p+1}_{i,-p} = \frac{i\psi_{i,0}(B^p_{i,-p}\psi_{i,-1} + A^p_{i,-p}\varphi_{i,-1})}{p} +\\
+ \frac{i\psi_{i,-1}(A^p_{i,-p}\varphi_{i,0} + B^p_{i,-p}\psi_{i,0} + A^p_{i,-p+1}\varphi_{i,-1} + B^p_{i,-p+1}\psi_{i,-1})}{p-1} - \frac{iA^p_{i,-p+1}(p-1)}{2},
\end{gathered}
\end{equation}
\begin{equation}
\begin{gathered}
B^{p+1}_{i,-p} = -\frac{i\varphi_{i,0}(B^p_{i,-p}\psi_{i,-1} + A^p_{i,-p}\varphi_{i,-1})}{p} -\\
 -\frac{i\varphi_{i,-1}(A^p_{i,-p}\varphi_{i,0} + B^p_{i,-p}\psi_{i,0} + A^p_{i,-p+1}\varphi_{i,-1} + B^p_{i,-p+1}\psi_{i,-1})}{p-1} + \frac{iB^p_{i,-p+1}(p-1)}{2}.\\
\end{gathered}
\end{equation}
Suppose that $p$ is odd number and $p=2k+1$. By (4), Lemma 1 and the induction assumption  it follows that
\begin{equation}
\frac{A^{2k+2}_{i,-2k-1}}{\psi_{i,-1}} + \frac{B^{2k+2}_{i,-2k-1}}{\varphi_{i,-1}}  = -ik\alpha_{2k+1}(\varphi_{i,0}\psi_{i,-1} + \psi_{i,0}\varphi_{i,-1})
\end{equation}
If  $p$ is even and $p=2k$, then
\begin{equation*}
\begin{gathered}
\frac{A^{2k+1}_{i,-2k}}{\psi_{i,-1}} - \frac{B^{2k+1}_{i,-2k}}{\varphi_{i,-1}}  =\\
(\varphi_{i,0}\psi_{i,-1} + \psi_{i,0}\varphi_{i,-1})\frac{i((8k-2)A^{2k}_{i,-2k} + \alpha_{2k}k\psi_{i,-1}(4\varphi_{i,-1}\psi_{i,-1} -(2k-1)^2)}{2(2k-1)k\psi_{i,-1}}=\\
=(\varphi_{i,0}\psi_{i,-1} + \psi_{i,0}\varphi_{i,-1})\frac{i((8k-2)\frac{A^{2k}_{i,-2k}}{\psi_{i,-1}} + \alpha_{2k}k(4n_i^2 -(2k-1)^2)}{2(2k-1)k}=\\
=(\varphi_{i,0}\psi_{i,-1} + \psi_{i,0}\varphi_{i,-1})\frac{i((8k-2)\beta\prod\limits^{l=k-1}_{l=1}(n_i^2 - l^2) + \alpha_{2k}k(4n_i^2 -(2k-1)^2)}{2(2k-1)k},
\end{gathered}
\end{equation*}
where $\beta$ doesn't depend on $\varphi_{i,-1}$, $\psi_{i,-1}$ and $n_i$. Easy to see that  $\alpha_l \neq 0$ for any $n_i \in \mathbb{N}$ and $l$. We know that there exists $l$ such that $f_{l}=0$ and $g_{l}=0$. So $A^l_{i,-l+1} = B^l_{i,-l+1}=0 \Rightarrow \frac{A^l_{i,-l+1}}{\varphi_{i,-1}} + (-1)^l \frac{B^l_{i,-l+1}}{\psi_{i,-1}} = 0$. But we see that $\alpha_l \neq 0$ and hence $\frac{A^l_{i,-l+1}}{\varphi_{i,-1}} + (-1)^l \frac{B^l_{i,-l+1}}{\psi_{i,-1}} =0$ if and only if $\varphi_{i,0}\psi_{i,-1} + \psi_{i,0}\varphi_{i,-1} = 0$. \\
\textbf{The Lemma is proved.}\\
\\
Let us assume that
 \begin{equation}
 \psi_{i,-1}\varphi_{i,p} + (-1)^p\varphi_{i,-1}\psi_{i,p}=0, \quad \frac{A^{t}_{i,p-t+1}}{\psi_{i,-1}} + (-1)^{p-t}\frac{B^{t}_{i,p-t+1}}{\varphi_{i,-1}} =0 ,
 \end{equation}
where $p=0,...,k-1$ and $t\in \mathbb{N}$ is arbitrary number.\\
Obviously
\begin{equation*}
\frac{A^{1}_{i,k}}{\psi_{i,-1}} + (-1)^{k+1}\frac{B^1_{i,k}}{\varphi_{i,-1}} = -i(\frac{\psi_{i,k}}{\psi_{i,-1}} + (-1)^k\frac{\varphi_{i,k}}{\varphi_{i,-1}}) = -\frac{i}{n^2}(\varphi_{i,-1}\psi_{i,k} + (-1)^k\psi_{i,k}\varphi_{i,-1}).
\end{equation*}
Consider
\begin{equation*}
\begin{gathered}
f_r = \frac{A^r_{i,-r}}{(x-a_i)^r} + ... + A^r_{i,0} + A^r_{i,1}(x-a_i) + .. + A^r_{i,k-r}(x-a_i)^{k-r} + O((x-a_i)^{k-r+1})\\
g_r = \frac{B^r_{i,-r}}{(x-a_i)^r} + ... + B^r_{i,0} + B^r_{i,1}(x-a_i) + .. + B^r_{i,k-r}(x-a_i)^{k-r} + O((x-a_i)^{k-r+1})
\end{gathered}
\end{equation*}
and suppose that
\begin{equation}
\frac{A^{r-1}_{i,k-r+2}}{\psi_{i,-1}} + (-1)^{k-r+1}\frac{B^{r-1}_{i,k-r+2}}{\varphi_{i,-1}} =   \alpha_{r-1, k-r+2}(\varphi_{i,-1}\psi_{i,k} + (-1)^k\psi_{i,k}\varphi_{i,-1}),
\end{equation}
where $\alpha_{r-1, k-r+2}$ doesn't depend on $C_i$ and coefficients $\varphi_{i,j}$, $\psi_{i,j}$. But $\alpha_{r-1, k-r+2}$ depends on $n_i^2 = \varphi_{i,-1}\psi_{i,-1}$.\\
Let us prove that
\begin{equation}
\frac{A^{r}_{i,k-r+1}}{\psi_{i,-1}} + (-1)^{k-r}\frac{B^{r}_{i,k-r+1}}{\varphi_{i,-1}} = \alpha_{r, k-r+1}(\varphi_{i,-1}\psi_{i,k} + (-1)^k\psi_{i,k}\varphi_{i,-1}).
\end{equation}
Denote
\begin{equation*}
pf_{r-1} + qg_{r-1} = \frac{h_{i,-r}}{(x-a_i)^{r}} + ... + h_{i,k-r+1}(x-a_i)^{k-r+1} + O((x-a_i)^{k-r+1}).
\end{equation*}
Let us consider two cases.\\
1) $k - r$ is even number. \\
Using (13), we get \\
\begin{equation*}
\begin{gathered}
h_{i,k-r}=\varphi_{i,-1}A^{r-1}_{i,k-r+1} +\psi_{i,-1}B^{r-1}_{i,k-r+1} +...+\varphi_{i,k-1}A^{r-1}_{i,-r+1} + \psi_{i,k-1}B^{r-1}_{i,-r+1} =\\
=\varphi_{i,-1}\psi_{i,-1}((\frac{A^{r-1}_{i,k-r+1}}{\psi_{i,-1}} + \frac{B^{r-1}_{i,k-r+1}}{\varphi_{i,-1}})+...+(\frac{\varphi_{i,k-1}A^{r-1}_{i,-r+1}}{\varphi_{i,-1}\psi_{i,-1}} + \frac{\psi_{i,k-1}B^{r-1}_{i,-r+1}}{\psi_{i,-1}\varphi_{i,-1}}))=0
\end{gathered}
\end{equation*}
Similarly, under the condition of (13), we have $h_{i,2t}=0$, where $2t \leqslant k-r$.
From (1) it follows that
\begin{equation*}
\frac{f_{r}}{\psi_{i,-1}} + \frac{g_{r}}{\varphi_{i,-1}} = \frac{i}{2}(\frac{f'_{r-1}}{\psi_{i,-1}} - \frac{g'_{r-1}}{\varphi_{i,-1}}) +i(\frac{p}{\varphi_{i,-1}}-\frac{q}{\psi_{i,-1}})(\int(pf_{r-1} + qg_{r-1})dx + C_{r-1}).
\end{equation*}
Note that
\begin{equation}
\int(pf_{r-1} + qg_{r-1})dx = \sum\limits^{t=k-r}_{t = -r} \frac{h_{i,t}(x-a_i)^{t+1}}{t+1} +  \frac{h_{i,k-r+1}(x-a_i)^{k-r+1}}{k-r} + O((x-a_i)^{k-r+1}),
\end{equation}
where $h_{i,p}=0$, $p$ is even number and $p\leqslant k-r$. We mentioned before that $f_k$ and $g_k$ are polynomials in $p,q,p',q',...,p^{(k-1},q^{(k-1)}$ for all $k$. Hence $h_{i,-1}=0$.\\
Moreover, we see from (13) that
\begin{equation}
\frac{p}{\varphi_{i,-1}}-\frac{q}{\psi_{i,-1}} = \sum\limits_{t=0}^{2t < k}(\frac{\varphi_{i,2t}}{\varphi_{i,-1}} - \frac{\psi_{i,2t}}{\psi_{i,-1}} )(x-a_i)^{2t} + (\frac{\varphi_{i,k}}{\varphi_{i,-1}} - \frac{\psi_{i,k}}{\psi_{i,-1}})(x-a_i)^{k} + ...
\end{equation}
Combining (16) and (17) and using the fact that $k-r+1$ is odd number, we obtain
\begin{equation}
\begin{gathered}
\frac{A^{r}_{i, k-r+1}}{\psi_{i,-1}} + \frac{B^{r}_{i, k-r+1}}{\varphi_{i,-1}} = \frac{i(k-r+2)}{2}(\frac{A^{r-1}_{i,k-r+2}}{\psi_{i,-1}} - \frac{B^{r-1}_{i, k-r+2}}{\varphi_{i,-1}})= \\
= \alpha_{r-1, k-r+2}\frac{i(k-r+2)}{2}(\varphi_{i,-1}\psi_{i,k} + (-1)^k\psi_{i,-1}\varphi_{i,-1}).
\end{gathered}
\end{equation}
So, if $k-r$ is even, then assumption (15) is true.\\
\\
2) Suppose that $k-r$ is odd number.\\
Arguing as before we see that
\begin{equation*}
\begin{gathered}
h_{i,k-r+1}=\varphi_{i,-1}A^{r-1}_{i,k-r+2} +\psi_{i,-1}B^{r-1}_{i,k-r+2} +...+\varphi_{i,k-1}A^{r-1}_{i,-r+1} + \psi_{i,k-1}B^{r-1}_{i,-r+1} =\\
=\varphi_{i,-1}\psi_{i,-1}((\frac{A^{r-1}_{i,k-r+2}}{\psi_{i,-1}} + \frac{B^{r-1}_{i,k-r+2}}{\varphi_{i,-1}})+...+(\frac{\varphi_{i,k-1}A^{r-1}_{i,-r+2}}{\varphi_{i,-1}\psi_{i,-1}} + \frac{\psi_{i,k-1}B^{r-1}_{i,-r+2}}{\psi_{i,-1}\varphi_{i,-1}})+\\
+(\frac{\varphi_{i,k}A^{r-1}_{i,-r+1}}{\varphi_{i,-1}\psi_{i,-1}} + \frac{\psi_{i,k}B^{r-1}_{i,-r+1}}{\psi_{i,-1}\varphi_{i,-1}}))=\psi_{i,-1}B^{r-1}_{i,-r+1}(\frac{\varphi_{i,k}}{\varphi_{i,-1}} + (-1)^{r-1}\frac{\psi_{i,k}}{\psi_{i,-1}})=\\
=\beta\prod\limits^{l=[\frac{r-1}{2}]}_{l=1}(n^2 - l^2)(\frac{\varphi_{i,k}}{\varphi_{i,-1}} + (-1)^{k}\frac{\psi_{i,k}}{\psi_{i,-1}}),
\end{gathered}
\end{equation*}
where $\beta$ is constant and doesn't depend on coefficients of series.\\
Easy to see from (13) that $h_{i,2t+1}=0$, where $2t + 1 < k-r+1$. From (1) it follows that
\begin{equation*}
\frac{f_{r}}{\psi_{i,-1}} - \frac{g_{r}}{\varphi_{i,-1}} = \frac{i}{2}(\frac{f'_{r-1}}{\psi_{i,-1}} + \frac{g'_{r-1}}{\varphi_{i,-1}}) -i(\frac{p}{\varphi_{i,-1}}+\frac{q}{\psi_{i,-1}})(\int(pf_{r-1} + qg_{r-1})dx + C_{r-1}).
\end{equation*}
From (13) we obtain that
\begin{equation*}
\begin{gathered}
\frac{p}{\varphi_{i,-1}}+\frac{q}{\psi_{i,-1}} = \sum\limits_{t=-1}^{t=k-1 }(\frac{\varphi_{i,t}}{\varphi_{i,-1}} + \frac{\psi_{i,t}}{\psi_{i,-1}} )(x-a_i)^{t} + (\frac{\varphi_{i,k}}{\varphi_{i,-1}} + \frac{\psi_{i,k}}{\psi_{i,-1}})(x-a_i)^{k} + ...=\\
= \frac{2}{(x-a_i)} + 2\sum\limits_{t=0}^{2t \leqslant k-2}\frac{\varphi_{i,2t+1}}{\varphi_{i,-1}}(x-a_i)^{2t+1} + (\frac{\varphi_{i,k}}{\varphi_{i,-1}} + \frac{\psi_{i,k}}{\psi_{i,-1}})(x-a_i)^{k}...
\end{gathered}
\end{equation*}
Using the fact that $k-r+1$ is even number, we get
\begin{equation*}
\begin{gathered}
\frac{A^{r}_{i, k-r+1}}{\psi_{i,-1}} - \frac{B^{r}_{i, k-r+1}}{\varphi_{i,-1}} = \frac{i(k-r+2)}{2}(\frac{A^{r-1}_{i,k-r+2}}{\psi_{i,-1}} + \frac{B^{r-1}_{i, k-r+2}}{\varphi_{i,-1}}) -2\frac{h_{i,k-r+1}}{k-r+1}=\\
= (\frac{i(k-r+2)\alpha_{r-1, k-r+2}}{2} - 2\beta\prod\limits^{l=[\frac{r-1}{2}]}_{l=1}(n^2 - l^2) )(\frac{\varphi_{i,k}}{\psi_{i,-1}} + (-1)^{k}\frac{\psi_{i,k}}{\varphi_{i,-1}}).
\end{gathered}
\end{equation*}
We obtain that if $k-r$ is odd, then assumption (15) is correct.\\
So, we proved that assumption (15) is correct for all $k$ and $r$.\\
\\
Suppose that $k<2n_i+1$. We know from Lemma 1 that $f_k \neq 0$ and $g_k\neq 0$ because  $A^k_{i,-k} \neq 0$.
We obtain that if there exists $m$ such that $f_{2m+1}=0$ and $g_{2m+1}=0$, then $A^{2m+1}_{i,p}=B^{2m+1}_{i,p}=0 \Rightarrow A^{2m+1}_{i,p} + (-1)^{p+1}B^{2m+1}_{i,p}=0$, where $p \leqslant 0$, $m\geqslant n_i$. Note that $A^{2n_i+1}_{i,-2n_i-1}=B^{2n_i+1}_{i,2n_i+1}=0$. \\
\\
We see from (18) that $\frac{A^{2n_i+2}_{i,-1}}{\psi_{i,-1}} + \frac{B^{2n_i+2}_{i,-1}}{\varphi_{i,-1}}=0$. We get from (15) that
\begin{equation*}
\frac{A^{2m+1}_{i,k-2m}}{\psi_{i,-1}} + (-1)^{k-2m+1}\frac{B^{2m+1}_{i,k-2m}}{\varphi_{i,-1}} =   \alpha_{2m+1, k-2m }(\frac{\varphi_{i,k}}{\psi_{i,-1}} + (-1)^{k}\frac{\psi_{i,k}}{\varphi_{i,-1}})=0.
\end{equation*}
Easy to see that $\alpha_{2m+1, k-2m}\neq 0$, where $k-2m<-1$. This means that $\frac{\varphi_{i,k}}{\psi_{i,-1}} + (-1)^{k}\frac{\psi_{i,k}}{\varphi_{i,-1}} = 0 \Rightarrow \varphi_{i,k}=(-1)^{k+1}\frac{\psi_{i,k}\varphi_{i,-1}^2}{n^2}$ for any $k \leqslant 2n_i-1$.\\
\\
\textbf{Theorem 1 is proved.}
\\
\begin{center}
\textbf{4. Proof of Theorem 2}
\end{center}

Let us consider the equation
\begin{equation}
M= \begin{pmatrix}
\imath & 0  \\
0 & -\imath  \\
\end{pmatrix}
\partial_xy +
 \begin{pmatrix}
0 & -\imath q(x)  \\
\imath p(x) & 0  \\
\end{pmatrix}y = zy
\end{equation}
where $y=(y_1(x),y_2(x))^{t}$. Solution of (19) have poles only at poles of $p$ and $q$. \\
Assume that
\begin{equation*}
\begin{gathered}
y_1=\alpha_0(x-a_i)^{\sigma} + \alpha_1(x-a_i)^{\sigma+1}+...\\
y_2=\beta_0(x-a_i)^{\sigma} + \beta_1(x-a_i)^{\sigma+1}+...
\end{gathered}
\end{equation*}
where $a_i$ is a pole of functions $p(x)$ and $q(x)$. Substituting $y$ in (19) and considering the coefficient of $(x-a_i)^{\sigma-1}$ we get
\begin{equation*}
 \begin{cases}
   \sigma\alpha_0- \psi_{i,-1}\beta_0=0\\
   \varphi_{i,-1}\alpha_0 -\sigma\beta_0  =0.\\
 \end{cases}
\end{equation*}
There exist solutions $\alpha_0, \beta_0 \neq 0$ if and only if $\sigma = \pm n_i$, where $n_i \in \mathbb{N}$.  Consider the coefficient of $(x-a_i)^{k+\sigma}$,  where $k$ is integer and $k\geqslant 0$. We obtain
\begin{equation*}
 \begin{cases}
   (k+\sigma)\alpha_{k-n_i+1}- \psi_{i,-1}\beta_{k-n_i+1}=F_1\\
   \varphi_{i,-1}\alpha_{k-n_i+1} -(k+\sigma)\beta_0 =F_2\\
 \end{cases}
\end{equation*}
where $F_1$ and $F_2$ don't depend  on $\alpha_{k-n_i+1}$ and $\beta_{k-n_i+1}$. Determinant of this system is equal to zero if and only if $k = -2\sigma$ and $k=0$. So, we have problems with finding coefficients only when $\sigma=-n_i$ and $k=2n_i$ because $k\geqslant 0$.\\
Let us prove that if $k\leqslant 2n_i-1$, then
\begin{equation}
n_i\beta_k + (-1)^k\varphi_{i,-1}\alpha_k =0
\end{equation}
or equivalently
\begin{equation*}
n_i\alpha_k + (-1)^k\psi_{i,-1}\beta_k=0.
\end{equation*}
We see that  $n_i\alpha_0 + \psi_{i,-1}\beta_0=0$ or equivalently $\varphi_{i,-1}\alpha_0 + n\beta_0=0$. Then
\begin{equation}
\begin{cases}
  (1-n_i)\alpha_1 - \psi_{i,-1}\beta_1  - \psi_{i,0}\beta_0=z\alpha_0\\
   (n_i-1)\beta_1 + \varphi_{i,-1}\alpha_1 + \varphi_{i,0}\alpha_0 =z\beta_0.\\
 \end{cases}
\end{equation}
Multiplying the first equation of (21) by $n_i$, the second by $\psi_{i,-1}$ and summing, we get
\begin{equation*}
n_i\alpha_{1} - \psi_{i,-1}\beta_1 + \psi_{i,-1}\varphi_{i,0}\alpha_0 - n\psi_{i,0}\beta_0=0
\end{equation*}
Under the condition of the Theorem we obtain that $n_i\alpha_{1} - \psi_{i,-1}\beta_1=0$. So, if $k=1$, then assumption (20) is true. Similarly, let us prove (20) for $k+1$ assuming that (20) is true for $k$, where $k< 2n_i-1$. Calculations show\\
\begin{equation}
\begin{cases}
  (n_i - k)\alpha_k - \psi_{i,-1}\beta_k - \psi_{i,0}\beta_{k-1}-...- \psi_{i,k-1}\beta_0=z\alpha_{k-1}\\
   -(n_i - k)\beta_k + \varphi_{i,-1}\alpha_k + \varphi_{i,0}\alpha_{k-1} + ...+\varphi_{i,k-1}\alpha_0 =z\beta_{k-1}.\\
 \end{cases}
\end{equation}
Let us multiply the first equation of (22) by $n_i$ the second by $(-1)^{k+1}\psi_{i,-1}$  and add. Easy to see that under the conditions of the Theorem we get (20).\\
Now let us consider coefficient of $(x-a_i)^{n_i-1}$
\begin{equation*}
\begin{cases}
n_i\alpha_{2n_i} - \psi_{i,-1}\beta_{2n_i} - \psi_{i,0}\beta_{2n_i-1} - ... -\psi_{2n_i-1}\beta_0 = z\alpha_{2n_i-1}\\
 \varphi_{i,-1}\alpha_{2n_i} + \varphi_{i,0}\alpha_{2n_i-1} + ... +\varphi_{2n_i-1}\alpha_0 - n\beta_{2n_i} = z\beta_{2n_i-1}\\
 \end{cases}
\end{equation*}
\begin{equation*}
\begin{cases}
n_i\alpha_{2n_i}  = z\alpha_{2n_i-1} + \psi_{i,-1}\beta_{2n_i} +  \psi_{i,0}\beta_{2n_i-1} + ... +\psi_{2n_i-1}\beta_0\\
  \varphi_{i,-1}\alpha_{2n_i} = n_i\beta_{2n_i} -  \varphi_{i,0}\alpha_{2n_i-1} - ... -\varphi_{2n_i-1}\alpha_0 + z\beta_{2n_i-1}.  \\
 \end{cases}
\end{equation*}
This system has solutions if and only if
\begin{equation*}
\begin{gathered}
z \frac{\alpha_{2n_i-1}}{n_i} + \frac{\psi_{i,-1}\beta_{2n_i}}{n_i} +  \frac{\psi_{i,0}\beta_{2n_i-1}}{n_i} + ... +\frac{\psi_{2n_i-1}\beta_0}{n_i} -\\ -n_i\frac{\beta_{2n_i}}{\varphi_{i,-1}} -  \frac{\varphi_{i,0}\alpha_{2n_i-1}}{\varphi_{i,-1}} - ... -\frac{\varphi_{2n_i-1}\alpha_0}{\varphi_{i,-1}} - z\frac{\beta_{2n_i-1}}{\varphi_{i,-1}}=0.\\
\end{gathered}
\end{equation*}
But we know from Lemma 1 that $\varphi_{i,-1}\psi_{i,-1}=n^2$ and under the conditions of the Theorem and (20) we have
\begin{equation*}
\begin{gathered}
\beta_{2n_i}(\frac{\varphi_{i,-1}\psi_{i,-1} -n^2_i}{n_i\varphi_{i,-1}}) + \frac{\varphi_{i,-1}\psi_{i,0}\beta_{2n_i-1} - n_i\varphi_{i,0}\alpha_{2n_i-1}}{n_i\varphi_{i,-1}} +...+ \frac{\varphi_{i,-1}\psi_{i,2n_i-1}\beta_{0} - n_i\varphi_{i,2n_i-1}\alpha_{0}}{n_i\varphi_{i,-1}} +\\ +z(\frac{\varphi_{i,-1}\alpha_{2n_i-1} - n_i\beta_{2n_i-1}}{n_i\varphi_{i,-1}}) = \\
=\varphi_{i,0}(\frac{\psi_{i,-1}\beta_{2n_i-1} - n_i\alpha_{2n_i-1}}{n_i\varphi_{i,-1}}) +...+ \varphi_{i,2n_i-1}\frac{-\psi_{i,-1}\beta_{0} -n_i\alpha_{0}}{n_i\varphi_{i,-1}} +\\ +z(\frac{\varphi_{i,-1}\alpha_{2n_i-1} - n_i\beta_{2n_i-1}}{n_i\varphi_{i,-1}})=0.
\end{gathered}
\end{equation*}
So, $y=(y_1(x),y_2(x))^{t}$, where $\sigma=\pm n_i$, are two linear independent formal solutions then it is also an actual solution near $a_i$ (see \cite{Levinson}, \S 4.3).
Finally, we obtain that solutions of (21) are locally meromorphic for all $z \in \mathbb{C}$.
\\
\textbf{Theorem is proved.}
\\
\\
\begin{center}
\textbf{5. Proof of Theorem 4}
\end{center}

Suppose $u(x)$ is a rational function  or a simply periodic function or an elliptic function. Suppose that $u(x)$ hasn't isolated singularity at infinity. Assume that $u(x)$  has finitely many poles in the  period strip or fundamental parallelogram. Suppose that function $u(x)$ has poles at points $a_1,a_2,..a_{N}$. In a neighborhood of $a_i$
\begin{equation*}
u(x) = \frac{a_{i,-2}}{(x-a_i)^2} + \frac{a_{i,-1}}{(x-a_i)} + a_{i,0} + a_{i,1}(x-a_i) + O((x-a_i)^2).
\end{equation*}
Theorem 3 follows from the following theorem ~\cite{weikard1}\\
\textbf{Theorem.}\\
\emph{Function $u(x)$ is Schrodinger  finite-gap potential if and only if
\begin{equation*}
a_{i,-2}=n_i(n_i+1), a_{i,-1}=0, a_{i,1}=0, ... , a_{2n_i - 1} = 0.
\end{equation*}}
But we have already proved that in a neighborhood of poles $a_i$ the Laurent series of function $\frac{n+1}{n}pq$  have the form
\begin{equation*}
\frac{n_i(n_i+1)}{(x-a_i)^2} + \sum^{t=n_i}_{t=0}h_{i,2t}(x-a_i)^{2t} + h_{2n_i+1}(x-a_i)^{2n_i+1} + O((x-a_i)^{2n_i+2})
\end{equation*}
\textbf{Theorem 3 is proved.}\\

Department of Geometry and Topology, Faculty of Mechanics and Mathematics, Lomonosov Moscow State University, Moscow, 119991 Russia.\\\\
E-mail address: vardan.o@mail.ru

\end{document}